OpenStudy (sam2211):

evaluate the following absolute value expressions (-100)-98=? 5(2-8)=? (-13)+(0)=? 14-(-10+3)

OpenStudy (polymorphism):

I assume the ( ) are suppose to be | | if so |-100|-98 100-98 =? 5(2-8)=? 2-8 = -6 |-6| = 6 5 * 6 = ? (-13)+(0)=? |-13| = 13 and 0=0 13 + 0 = ?

OpenStudy (sam2211):

yeah

OpenStudy (polymorphism):

14-(-10+3) do the work in the () || symbols 14-(-10+3) |-10+3| = |-7| = 7 14 - 7 = ?

OpenStudy (sam2211):

I just don't understand

OpenStudy (polymorphism):

Absolute value is a distance measurement. It does not matter if it is negative or positive. So if we have a negative such as |-7| this equals 7 and |7| = 7 So when we have something like 5|2-8|=? First we do the math in || symbols |2-8| |2-8|= |-6| = 6 So now we have 5 * 6 = 30

OpenStudy (sam2211):

OH I GET IT NOW THANK YOU!!